OpenStudy (anonymous):

Simplify (x2 + 7x +12)/(x2 -x -12)

OpenStudy (anonymous):

\[\frac{ x ^{2} + 7x + 12 }{ x ^{2} -x -12 } \] You use Bhaskara formula and gets that \[x ^{2} + 7x +12 = (x+4)(x+3)\] and do the same on \[x ^{2} -x - 12 = (x+3)(x-4)\] making \[\frac{ (x+4)(x+3) }{ (x-4)(x+3) }\] you get \[\frac{ (x+4) }{ (x-4) }\]

OpenStudy (anonymous):

Sorry, you might know it not by bhaskara formula but as quadratic formula